OpenStudy (anonymous):

A theater has 37 seats in the first row, 41 seats in the second row, 45 seats in the third row and so on. write the general term for a sequence (an) that give the number of seats in row n.

OpenStudy (raden):

this is a arithmatic series with the first term is 37, 2nd 41, ... therefore use formula : --------------- an = a1 + (n-1)d --------------- with : a1 = the first term d = different the 2nd and 1st term (41-37=4) so, just put a1= 37 and d=4 into that formula

OpenStudy (anonymous):

ok, and the question also asks how many seats are in row 17, I got 101 is that correct

OpenStudy (raden):

a17 = a1 + (17-1)*4 a17 = 37 + 16*4 a17 = 37 + 64 a17 = 101 u are correct

OpenStudy (anonymous):

ok. thank you very much

OpenStudy (raden):

welcome